OpenStudy (anonymous):

simplify by removing a factor equal to 1 50a^2b/40ab^3 need help

OpenStudy (anonymous):

??/ you need to add parenthesis

OpenStudy (anonymous):

is it a^2 or a^2b ?

OpenStudy (anonymous):

(5/4) * (a/b^2)

OpenStudy (anonymous):

50/40 reduces to 5/4

OpenStudy (anonymous):

a^2/a reduces to a/1

OpenStudy (anonymous):

b/b^3 reduces to 1/b^2

OpenStudy (anonymous):

remember that any number over the same number = 1 in other words, a/a = 1; b/b = 1 or even if you have a^2 / a, this is equivalent to a * a/a since a/a =1, you are left with a * 1

Directrix (directrix):